OpenStudy (anonymous):

What is the area of a circle with a circumference of 4π^2 ?

OpenStudy (jhannybean):

\[C=2\pi r~ , ~ C = 4\pi^2\]\[4\pi^2 = 2\pi r\]\[\color{red}{r} = \frac{4\pi^2}{2\pi} = ~?\]\[A=\pi \color{red}{r}^2\]